40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 82     Entry time: Thu Nov 8 00:55:44 2007
Author: pkp 
Type: Update 
Category: OMC 
Subject: Suspension tests 
[Sam , Pinkesh]

We tried to measure the transfer functions of the 6 degrees of freedom in the OMS SUS. To our chagrin, we found that it was very hard to get the OSEMs to center and get a mean value of around 6000 counts. Somehow the left and top OSEMs were coupled and we tried to see if any of the OSEMs/suspension parts were touching each other. But there is still a significant coupling between the various OSEMs. In theory, the only OSEMS that are supposed to couple are [SIDE] , [LEFT, RIGHT] , [TOP1, TOP2 , TOP3] , since the motion along these 3 sets is orthogonal to the other sets. Thus an excitation along any one OSEM in a set should only couple with another OSEM in the same same set and not with the others. The graphs below were obtained by driving all the OSEMS one by one at 7 Hz and at 500 counts ( I still have to figure out how much that is in units of length). These graphs show that there is some sort of contact somewhere. I cant locate any physical contact at this point, although TOP2 is suspicious and we moved it a bit, but it seems to be hanging free now. This can also be caused by the stiff wire with the peek on it. This wire is very stiff and it can transmit motion from one degree of freedom to another quite easily. I also have a graph showing the transfer function of the longitudnal degree of freedom. I decided to do this first because it was simple and I had to only deal with SIDE, which seems to be decoupled from the other DOFs. This graph is similar to one Norna has for the longitudnal DOF transfer function, with the addition of a peak around 1.8 Hz. This I reckon could very be due to the wire, although it is hard to claim for certain. I am going to stop the measurement at this time and start a fresh high resolution spectrum and leave it running over night.

There is an extra peak in the high res spectrum that is disturbing.
Attachment 1: shakeleft.pdf  5 kB  | Hide | Hide all
shakeleft.pdf
Attachment 2: shakeright.pdf  5 kB  | Hide | Hide all
shakeright.pdf
Attachment 3: shakeside.pdf  5 kB  | Hide | Hide all
shakeside.pdf
Attachment 4: shaketop1.pdf  5 kB  | Hide | Hide all
shaketop1.pdf
Attachment 5: shaketop2.pdf  5 kB  | Hide | Hide all
shaketop2.pdf
Attachment 6: shaketop3.pdf  5 kB  | Hide | Hide all
shaketop3.pdf
Attachment 7: LongTransfer.pdf  14 kB  | Hide | Hide all
LongTransfer.pdf LongTransfer.pdf LongTransfer.pdf
Attachment 8: Shakeleft7Nov2007_2.pdf  9 kB  Uploaded Thu Nov 8 11:22:37 2007  | Hide | Hide all
Shakeleft7Nov2007_2.pdf
Attachment 9: Shakeleft7Nov2007_2.png  25 kB  Uploaded Thu Nov 8 11:23:20 2007  | Hide | Hide all
Shakeleft7Nov2007_2.png
ELOG V3.1.3-